OpenStudy (anonymous):

If the area of a circle whose diameter is pi, is written as a pi to the b squared, what is the value of ab?

OpenStudy (anonymous):

where is the pic

OpenStudy (anonymous):

there is none

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Yes, I have skype

OpenStudy (anonymous):

no numbers

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

you don't any

OpenStudy (anonymous):

lookit

OpenStudy (anonymous):

pi/2

OpenStudy (anonymous):

really hmmmmmmmmm

OpenStudy (anonymous):

pi cubed/4

OpenStudy (anonymous):

that's as far as I have gotten

OpenStudy (anonymous):

the answer solution is 3/4

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

hmm i got to where u got also

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

can anyone help Mr. Noctowl here?

OpenStudy (anonymous):

i love pokemon

OpenStudy (anonymous):

The tcg format just rotated here in the U.S.

OpenStudy (anonymous):

Its now BW-on

OpenStudy (anonymous):

i no i watch in in japanese

OpenStudy (anonymous):

Nice

OpenStudy (anonymous):

Now

OpenStudy (anonymous):

Back to the problem

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

and then we can hold our discussion about pokemon

OpenStudy (anonymous):

what do you say, Aripotta?

OpenStudy (anonymous):

no i am helpping him

OpenStudy (aripotta):

i have no idea. i don't even know where a came into play here.

OpenStudy (anonymous):

no helping for u

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

krot

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

pi cubed/4

OpenStudy (anonymous):

im working it out

OpenStudy (aripotta):

pi cubed/4 is the area?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I don't know what a or b equals though

OpenStudy (anonymous):

I am assuming a equals 3 and b equals 4

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

mabye

OpenStudy (anonymous):

mabye b equals 1/3

OpenStudy (anonymous):

and a equals 4

OpenStudy (anonymous):

ehh?

OpenStudy (aripotta):

is ab the diameter?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

would the lenght of ab be twice

OpenStudy (anonymous):

the solution is 3/4

OpenStudy (anonymous):

the diameter is pi, actually

OpenStudy (anonymous):

the radius is pi/2

OpenStudy (anonymous):

wiat so what am i looking for

OpenStudy (anonymous):

the area is pi cubed/4

OpenStudy (anonymous):

well, they rewrote as a pi to the cubed power

OpenStudy (anonymous):

I mean

OpenStudy (anonymous):

a pi to the b power

OpenStudy (anonymous):

I am trying to find the value of ab

OpenStudy (anonymous):

ohhhh

OpenStudy (aripotta):

if ab is the diameter, then we already know it. it's pi.

OpenStudy (anonymous):

Son Goku?

hartnn (hartnn):

diameter = pi, radius = pi/2 AREA = pi / 4 = a pi b^2. so ab^2 = 1/4

OpenStudy (aripotta):

wat

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

pi cubed/4

OpenStudy (anonymous):

2 pi r

OpenStudy (anonymous):

I mean

OpenStudy (anonymous):

pi r squared

OpenStudy (anonymous):

from pi/2 we should have gone to pi cubed/4

OpenStudy (anonymous):

when we went from radius to area

OpenStudy (anonymous):

it will equal 1/4

OpenStudy (anonymous):

the answer solution is 3/4

OpenStudy (anonymous):

ab = 1/4

OpenStudy (anonymous):

I think I might know why

OpenStudy (anonymous):

/4 is the divisor

OpenStudy (anonymous):

so

OpenStudy (anonymous):

I am assuming that b becomes 1/4

OpenStudy (anonymous):

thats how we got from a cubed/4 to a pi to the b power

OpenStudy (anonymous):

a equals 3 from the cube root of pi

OpenStudy (anonymous):

so

OpenStudy (anonymous):

a=3

OpenStudy (anonymous):

b=1/4

OpenStudy (aripotta):

just gonna sit this one out..

OpenStudy (anonymous):

3 times 1/4=3/4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

srry i have to make dinner

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

good answer

OpenStudy (anonymous):

i guessing your right

OpenStudy (anonymous):

aripotta, son goku?

OpenStudy (anonymous):

goku

OpenStudy (anonymous):

son goku in Japanese and manga

OpenStudy (anonymous):

no it is martha

OpenStudy (anonymous):

lolz!

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

thats from dragon crystals

OpenStudy (anonymous):

or something

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

roy and marth